Heath wants 10 L of a solution that is 90% bleach. He will combine pure bleach with a mixture that is 20% bleach. Solve the system. How many liters of pure bleach will heath use. How many liters of the 20% mixture will he use
Answer by Alan3354(69443)

Heath wants 10 L of a solution that is 90% bleach. He will combine pure bleach with a mixture that is 20% bleach. Solve the system. How many liters of pure bleach will heath use. How many liters of the 20% mixture will he use
------------
p = pure bleach
t = 20% bleach
---
p + t = 10 (total solution)
100p + 20t = 10*90 (total bleach)
----
etc